THE NEGLECTED HALOGEN DIMER BANDS: DIMERIZATION EQUILIBRIUM CONSTANTS FROM SPECTROPHOTOMETRIC DATA

J. TELLINGHUISEN, Department of Chemistry, Vanderbilt University, Nashville, TN 37235.

The halogens exhibit concentration-dependent absorption bands in the UV that were long ago attributed to the formation of dimers. Although the spectral profiles of these bands have been moderately well characterized, the equilibrium constants K_{c} for their formation have mostly remained ill-defined. In particular, spectrophotometric data have been dismissed as inherently incapable of providing estimates of K_{c}. Here this presumption is shown to be incorrect in the cases of Br_{2}(g) dimer absorption near 205 nm and I_{2}(soln) absorption near 300 nm. The analysis is enabled by precise spectrophotometric data for 20 concentrations, analyzed via a multiwavelength global approach. The Br_{2} dimerization constant is found to be K_{c} = 2.5(3) L/mol at 22C, giving less than 2 percent dimerization at the maximum studied pressure (119 Torr). The K_{c} value is more than a factor of 2 larger than a previous estimate from analysis of PVT data – an approach that can be shown to yield at best a lower limit.